A teacher conducted a study to know if blended learning affects the students’ performances. A class of 30 students of Grade 11 was surveyed and found out that their mean score was 83 with a standard deviation of 4. A study from other country revealed that with a standard deviation of 3. Test the hypothesis at 0.10 level of significance.

Parameter: _______________________________________

Statistic: ________________________________

Null Hypothesis: ____________________________________________________

Alternative hypothesis: _____________________________________________


1
Expert's answer
2022-05-08T14:14:48-0400

Parameter: Difference of two independent normal variables "\mu_{X-Y}"

Let "X" have a normal distribution with mean "\mu_X" and variance "\sigma_X^2."

Let "Y" have a normal distribution with mean "\mu_Y" and variance "\sigma_Y^2."

If "X" and "Y"are independent, then "X-Y"will follow a normal distribution with mean "\mu_X-\mu_Y" and variance "\sigma_X^2+\sigma_Y^2."

"\mu_{X-Y}=83-80=3""s_{X-Y}=\sqrt{(3)^2+(4)^2}=5"



Statistic: "t-" statistic


The following null and alternative hypotheses need to be tested:

"H_0:\mu=0"

"H_1:\mu\not=0"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.10,"

"df=n-1=30-1=29" degrees of freedom, and the critical value for a two-tailed test is "t_c=1.699127."

The rejection region for this two-tailed test is "R=\{t:|t|>1.699127\}"

The t-statistic is computed as follows:


"t=\dfrac{\mu_{X-Y}-\mu}{s_{X-Y}/\sqrt{n}}=\dfrac{3-0}{5/\sqrt{30}}\approx3.2863"

Since it is observed that "|t|=3.2863>1.699127=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed "\alpha=0.10," "df=29, t=3.2863" is "p=0.002659," and since "p=0.002659<0.10=\alpha," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 0, at the "\alpha=0.10" significance level.

Therefore, there is enough evidence to claim that blended leaming affects the students' performances, at the "\alpha=0.10" significance level.